# MToP: A MATLAB Benchmarking Platform for Evolutionary Multitasking

### v5

@@ 3.3.3. Performance Metric
Unlike the implementation of algorithm and problem classes, metric codes in MToP are defined as functions. An illustrative example of metric implementation is presented in Listing LABEL:list:metric. The input parameter of the function is MTOData, which is generated during experimental execution. The function returns result, comprising properties such as Metric, RowName, ColumnName, TableData, ConvergeData, and ParetoData. The Metric property can take values of either Min or Max, indicating whether a smaller or larger metric value is preferable. RowName, ColumnName, and TableData are utilized to present metric results in the GUI table. On the other hand, ConvergeData and ParetoData are employed to exhibit metric convergence results for convergence plots and non-dominated solutions for Pareto plots, respectively. Note that specific metric calculation functions are not elaborated here for the sake of simplicity. The metrics currently integrated into MToP are enumerated in Table LABEL:tab:metric of the supplementary files.
@@ 4.1.3. Visualizing Algorithm Behavior
After configuring the parameters and starting the run, MToP offers dynamic display utilities such as Draw Obj and Draw Dec to enable researchers to explore algorithm behavior in real-time. Figure 9 illustrates an example of population variation of MO-MFEA on CEC17-MTMO4, and compares its final state with that of MTDE-MKTA, showcasing the evolution in both objective space and decision space. CEC17-MTMO4 is a multi-objective problem with two tasks, each comprising two objectives. The optimal Pareto set of the first task contains diversity dimension $1$ with optimal value range $[0,1]$ and convergence dimension $[2,50]$ with optimal value $0.5$. The optimal Pareto set of the second task contains diversity dimension $1$ with optimal value range $[0,1]$, convergence dimension $[2,40]$ with optimal value $0.5$, and convergence dimension $[41,50]$ with optimal value $0.5005$. Through these dynamic visualizations, researchers can intuitively grasp the algorithm’s operational behavior and delve deeper into its search dynamics.
@@
Figure 9. Schematic of population snapshots on CEC17-MTMO4 across generations. The top row (a)-(f) shows the objective space, where the blue dots are the algorithm’s non-dominated solutions and the black line is the true Pareto Frontier (PF). The bottom row (g)-(l) shows the decision space, where each black polyline represents the 50-dimensional decision vector of one individual in the population. Panels (a)-(e) and (g)-(k) depict MO-MFEA at Gen=1, 200, 500, 700, and 1000. Panels (f) and (l) show the example of MTDE-MKTA at Gen=1000.
In this example, during the initial phase of evolution (Gen=1), the population initialization is dispersed across the decision space and poorly situated in the objective space. As evolution progresses through the first and middle phases (Gen=200), individual gene knowledge transfer leads to populations swiftly converging to more favorable positions in the objective space, albeit at the expense of reduced diversity. Subsequently, in the middle and late stages (Gen=500 to 700), the population gradually emphasizes diversity. By Gen=1000, MO-MFEA achieves the optimal Pareto frontier on the first task but becomes trapped in local optima on the second task. The schematics provide the key insight: the population for the second task is clustered in the similar decision space region as the population for the first task. However, the true optimal regions of these two tasks do not precisely overlap in the decision space. This observation strongly suggests the presence of negative knowledge transfer: the strong convergence of Task 1 has incorrectly pulled the Task 2 population into its own optimal region, causing it to converge on a local optimum for Task 2. This visualization functionality in MToP is crucial for diagnosing such algorithmic failures.
Furthermore, MToP can be used to validate the effectiveness of remedial approaches. For instance, Figure 9 (f) and (l) show the final population snapshot of MTDE-MKTA, an algorithm with multiple knowledge transfer mechanisms designed to mitigate this issue. In contrast to MO-MFEA, MTDE-MKTA successfully converges to the true optimal Pareto frontier for both tasks. This is further confirmed in the decision space, where the inset plot clearly shows the convergence to the correct optimal values ($0.5$ for dimensions $[2,40]$ and $0.5005$ for $[41,50]$), thus overcoming the negative transfer problem. In MToP, the decision space population function applies not only to multi-objective problems but also to single-objective problems.
@@ 4.2.1. Setting Experiment Parameters
Experiment parameters are set in the Set Exp Parameter panel (Step 1). This includes setting the number of Repetitions (independent runs), the Rand Seed for reproducibility, and the Data Length (the number of convergence points to save per run). Toggles for Parallel execution and SaveDec (to save decision variables) are also located here. Enabling Parallel allows independent runs to execute simultaneously. The number of concurrent processes depends on the system’s processor cores, and MATLAB automatically handles the scheduling. The typical execution flow processes one problem at a time, running all added algorithms on it in parallel before moving to the next problem.
@@ 4.2.2. Selecting Algorithms and Problems
To assist researchers in this selection, especially when starting a new study, Table 5 provides a curated guide to benchmark suites and high-performing algorithms. The listed benchmarks are widely used and recognized within the EMT community, serving as standard testbeds. The recommended algorithms are not arbitrary; they represent the top 5 algorithms based on the Friedman ranking of their average metric performance over 30 independent runs on each benchmark set. The full experimental validation for this ranking is detailed later in Section 5.1. This table serves as an excellent starting point for researchers and can also offer valuable insights when selecting algorithms for new real-world problems. Comprehensive lists, references, and detailed descriptions of all problems, algorithms, and metrics available in MToP are provided in Table LABEL:tab:prob-mt, Table LABEL:tab:algo-mt, and Table LABEL:tab:metric, respectively, in the supplementary files.
@@ 4.2.3. Running Experiment and Calculating Metrics
  • For standard multi-task problems (e.g., CEC17-MTSO, CEC19-MTMO), the number of tasks is small, making it practical and recommended to analyze per-task metrics directly, such as Obj, IGD+, or HV. For multi-objective metrics, IGD+ is generally recommended over the traditional IGD as it provides a more comprehensive and stable measure of both convergence and distribution. Furthermore, the HV metric is also provided, as it is a valuable indicator that does not require a known true Pareto front, making it suitable for real-world problems where the optimum is unknown.
@@ 4.2.5. Exporting Experiment Data and Results
  • IOHanalyzer Data (csv): This option exports the complete convergence history (not just the final values) for all selected runs into a csv format. This format is specifically structured for use with the IOHanalyzer (Wang2022IOHanalyzer), facilitating advanced external analysis.
  • Best Decision Variable (mat): This option extracts the decision variables of the best-found solutions of all algorithms and repetitions and saves them to a .mat file for further inspection or use in other applications. For single-objective MTO problems, the best solution is determined by the lowest objective value. For multi-objective MTO problems, the best solutions are a set of Pareto sets (Liu2024Finding; Choong2022Jack).
@@ 4.2.6. Visualizing Metric Convergence and Pareto Front
Finally, the module offers integrated plotting tools at the bottom of the results panel (Step 6). By selecting one or more rows and columns in the results table, the user can generate plots. Clicking the Converge button generates a convergence plot for the selected metric results, with a Log toggle available to switch the y-axis to a logarithmic scale, and a Range toggle to display the 95% confidence interval around the mean curve. For multi-objective problems, clicking the Pareto button will plot the final acquired non-dominated solutions (Pareto fronts). Additionally, the convergence data can be exported in a format compatible with IOHanalyzer (Wang2022IOHanalyzer) for further external analysis and visualization.
@@ 4.4. Executing via Alternative Command Line
A comprehensive set of examples demonstrating this full process including how to instantiate and modify algorithm/problem objects, run experiments, calculate metrics like Obj and IGD, and programmatically plot the final convergence curves and Pareto fronts is provided in the cmd_examples.m script located in MToP’s root directory.
@@ 5. Validation and Comparison
This section first performs reproducible validation tests to ensure that the algorithms and problems implemented in MToP are reliable and accurate. Following this, a performance comparison between MToP and a popular EC platform, PlatEMO (Tian2017PlatEMO), is conducted to demonstrate MToP’s comparable efficiency.
@@ 5.2. Comparison with Other Platforms
Beyond internal validation, it is crucial to benchmark MToP’s performance and efficiency against established, state-of-the-art platforms. We selected PlatEMO (Tian2017PlatEMO), a widely-used and highly-regarded platform for evolutionary computation, as the basis for our comparison. To ensure a fair and direct comparison, we selected four popular and well-understood single-task multi-objective algorithms that are implemented in both platforms: NSGA-II, SPEA2, MOEA-D, and SMS-EMOA. We executed these algorithms on the standard DTLZ1-7 benchmark suite. Both platforms were run under identical experimental settings, including 30 independent runs and the use of parallel execution.
First, we compared the computational efficiency. Figure 12 displays the average wall-clock run time for each algorithm across the entire DTLZ suite. The results show that the execution times for MToP and PlatEMO are highly comparable. For NSGA-II, SPEA2, and MOEA-D, the run times are nearly identical. For SMS-EMOA, which involves computationally intensive hypervolume calculations, both platforms show similar overhead. This demonstrates that MToP’s architecture is lightweight and efficient, with no significant computational overhead compared to PlatEMO.
Second, we validated the correctness of our algorithmic implementations by comparing the final solution quality. Table LABEL:tab:results-dtlz presents the mean and standard deviation of the IGD metric for 30 runs. A Wilcoxon rank-sum test (at a 0.05 significance level) was conducted between the MToP and PlatEMO results for each case. The + / - / = summary shows that there is no statistically significant difference in the vast majority of cases (23 out of 28). The few minor differences are likely due to small implementation-level variations, but the overall performance is statistically indistinguishable. This similarity in performance is further confirmed visually in Figure LABEL:fig:pareto-comparison, which plots the final Pareto fronts from a sample run for each algorithm. The acquired fronts from both platforms are visually coincident.
@@ 6. Discussion and Outlook
Moving forward, the development and enhancement of MToP will continue, drawing inspiration and contributions from the academic community. It is our aspiration that MToP evolves into a valuable tool to propel research in the field of EMT and, by extension, advance the broader evolutionary computation field. Specifically, MToP is considering the incorporation of many-objective optimization codes (Cheng2016RVEA; Deb2014NSGA3), multitask combinatorial optimization codes (Feng2021VRPHTO; Feng2021EEMTA) and more real-world applications (Gupta2021Half; Huang2023Evolutionary) in the future.
@@ Acknowledgements.
This work was partly supported by the National Natural Science Foundation of China (Grant No. 62076225), the Major Research & Development Program of Hubei Province (Grant No. 2025BEB002 & 2023BCA006), the Regional Innovation System Program of Hubei Province (Grant No. 2025EIA022), and the Hubei Superior and Distinctive Discipline Group of “New-Energy Vehicles and Smart Transportation”. All algorithm, problem, and metric implementations within MToP adhere closely to the methodologies outlined in their respective articles. The code provided by the original authors undergoes modifications based on the source code during its integration into MToP. The authors would like to thank the researchers who assisted in modifying the code into the MToP pattern and the researchers who open-sourced their codes. Some of the function implementations in MToP about multi-objective optimization are referred to the code implementation in PlatEMO (Tian2017PlatEMO), for which the authors are also thankful.

### v6

@@ 3.3.3. Performance Metric
Unlike the implementation of algorithm and problem classes, metric codes in MToP are defined as functions. An illustrative example of metric implementation is presented in Listing LABEL:list:metric. The input parameter of the function is MTOData, which is generated during experimental execution. The function returns result, comprising properties such as Metric, RowName, ColumnName, TableData, ConvergeData, and ParetoData. The Metric property can take values of either Min or Max, indicating whether a smaller or larger metric value is preferable. RowName, ColumnName, and TableData are utilized to present metric results in the GUI table. On the other hand, ConvergeData and ParetoData are employed to exhibit metric convergence results for convergence plots and non-dominated solutions for Pareto plots, respectively. Note that specific metric calculation functions are not elaborated here for the sake of simplicity. The metrics currently integrated into MToP are enumerated in Table 5.
Table 5. Metrics in MToP. Abbreviations: SO/MO = single-/multi-objective; ST/MT = single-/multi-task; CV = constraint violation; FR = feasible rate; HV = hypervolume; IGD = inverted generational distance; IGD+ = improved IGD plus; AV/UV = average/unified average; MTS = multitask score; CMT = competitive multitask; NBR = number of best results.
Metric Objective Task Description
Obj SO ST/MT Objective value for each task
Obj (AV) SO MT Average Obj for all tasks
Obj (UV) SO MT Unified average Obj for all tasks
Obj (MTS) SO MT Multitask score of Obj for all tasks
Obj (CMT) SO MT Competitive multitask Obj for all tasks
Obj (NBR) SO MT Number of best Obj result for all tasks
CV SO ST/MT Constraint violation for each task
FR SO ST/MT Feasible rate for each task
HV MO ST/MT Hypervolume for each task
HV (MTS) MO MT Multitask score of HV for all tasks
HV (CMT) MO MT Competitive multitask HV for all tasks
IGD MO ST/MT Inverted generational distance for each task
IGD (AV) MO MT Average IGD for all tasks
IGD (MTS) MO MT Multitask score of IGD for all tasks
IGD (CMT) MO MT Competitive multitask IGD for all tasks
IGD+ MO ST/MT Improved IGD plus for each task
IGD+ (MTS) MO MT Multitask score of IGD+ for all tasks
IGD+ (CMT) MO MT Competitive multitask IGD+ for all tasks
Spread MO ST/MT Spread metric for each task
Spread (CMT) MO MT Competitive multitask spread for all tasks
Run Time SO/MO ST/MT Algorithm running time for all tasks
@@ 4.1.3. Visualizing Algorithm Behavior
After configuring the parameters and starting the run, MToP offers dynamic display utilities such as Draw Obj and Draw Dec to enable researchers to explore algorithm behavior in real-time. Fig. 9 illustrates an example of population variation of MO-MFEA (Gupta et al., 2017) on CEC17-MTMO4, and compares its final state with that of MTDE-MKTA (Li and Gong, 2025), showcasing the evolution in both objective space and decision space. CEC17-MTMO4 is a multi-objective problem with two tasks, each comprising two objectives. The optimal Pareto set of the first task contains diversity dimension $1$ with optimal value range $[0,1]$ and convergence dimension $[2,50]$ with optimal value $0.5$. The optimal Pareto set of the second task contains diversity dimension $1$ with optimal value range $[0,1]$, convergence dimension $[2,40]$ with optimal value $0.5$, and convergence dimension $[41,50]$ with optimal value $0.5005$.
Before analyzing the visual patterns, it is essential to formally distinguish the knowledge transfer mechanisms of the compared algorithms to understand the underlying causes of their behaviors. MO-MFEA relies on implicit knowledge transfer via assortative mating. Let $\bm{x}_{i}$ and $\bm{x}_{j}$ be parents from different tasks. The offspring $\bm{u}_{i}$ is generated using the simulated binary crossover, which directly mixes the decision variables:
(3) $\bm{u}_{i,d}=\begin{cases}0.5[(1+\gamma)\bm{x}_{i,d}+(1-\gamma)\bm{x}_{j,d}]&\text{if }r\leq 0.5\\ 0.5[(1-\gamma)\bm{x}_{i,d}+(1+\gamma)\bm{x}_{j,d}]&\text{otherwise}\end{cases}$
where $\gamma$ is the spread factor and $r$ is a random number. This mechanism lacks explicit domain adaptation, which may lead to negative transfer if the optimal regions of the tasks are misaligned.
In contrast, MTDE-MKTA employs an explicit knowledge transfer strategy based on evolution path maintenance. It models the population distributions of the source and target tasks as Gaussian distributions $\mathcal{N}(\bm{\mu}_{s},\bm{\sigma}_{s})$ and $\mathcal{N}(\bm{\mu}_{t},\bm{\sigma}_{t})$, respectively. A solution $\bm{x}$ from the source task is explicitly transformed into a candidate solution $\bm{y}$ for the target task by aligning their statistical distributions:
(4) $\bm{y}=\bm{\mu}_{t}+\frac{\bm{\sigma}_{t}\circ(\bm{x}-\bm{\mu}_{s})}{\bm{\sigma}_{s}}$
where $\circ$ denotes element-wise multiplication and division. This transformation adapts the transferred solution to the target task’s search space, thereby correcting the decision variable bias and mitigating negative transfer.
@@
Figure 9. Schematic of population snapshots on CEC17-MTMO4 across generations. The top row (a)-(f) shows the objective space, where the blue dots are the algorithm’s non-dominated solutions and the black line is the true Pareto Front. The bottom row (g)-(l) shows the 50-dimensional decision space. The horizontal axis represents the dimension index (1 to 50), and the vertical axis represents the unified decision variable value. Each individual in the population is visualized as a polyline connecting its values across all dimensions. The dense overlap of these polylines (appearing as black bands or lines) visually reflects the convergence consistency and the value distribution of the entire population. Panels (a)-(e) and (g)-(k) depict MO-MFEA at Gen=1, 200, 500, 700, and 1000. Panels (f) and (l) show the example of MTDE-MKTA at Gen=1000.
Through these dynamic visualizations, researchers can intuitively grasp the algorithm’s operational behavior. In this example, during the initial phase of evolution (Gen=1), the population initialization is dispersed across the decision space and poorly situated in the objective space. As evolution progresses through the first and middle phases (Gen=200), individual gene knowledge transfer leads to populations swiftly converging to more favorable positions in the objective space, albeit at the expense of reduced diversity. Subsequently, in the middle and late stages (Gen=500 to 700), the population gradually emphasizes diversity. By Gen=1000, MO-MFEA achieves the optimal Pareto front on the first task but becomes trapped in local optima on the second task. The schematics provide the key insight: the population for the second task is clustered in the similar decision space region as the population for the first task. However, the true optimal regions of these two tasks do not precisely overlap in the decision space. This observation strongly suggests the presence of negative knowledge transfer: the strong convergence of Task 1 has incorrectly pulled the Task 2 population into its own optimal region. This confirms the limitation of the implicit transfer mechanism (Eq. (3)) discussed above: without domain adaptation, the direct mixing of variables drags the population towards the wrong attractor.
Furthermore, MToP can be used to validate the effectiveness of remedial approaches. For instance, Fig. 9 (f) and (l) show the final population snapshot of MTDE-MKTA, an algorithm with multiple knowledge transfer mechanisms designed to mitigate this issue. In contrast to MO-MFEA, MTDE-MKTA successfully converges to the true optimal Pareto front for both tasks. This validates the efficacy of the explicit transformation (Eq. (4)), which successfully re-mapped the guiding solutions to the correct decision space region ($0.5005$), thus overcoming the negative transfer problem.
@@ 4.2.1. Setting Experiment Parameters
Experiment parameters are set in the Set Exp Parameter panel (Step 1). This includes setting the number of Repetitions (independent runs), the Rand Seed for reproducibility, and the Data Length (the number of checkpoints to save per run). Toggles for Parallel execution and SaveDec (to save decision variables) are also located here. Enabling Parallel allows independent runs to execute simultaneously. The number of concurrent processes depends on the system’s processor cores, and MATLAB automatically handles the scheduling. The typical execution flow processes one problem at a time, running all added algorithms on it in parallel before moving to the next problem.
@@ 4.2.2. Selecting Algorithms and Problems
To assist researchers in this selection, especially when starting a new study, Table 6 provides a curated guide to benchmark suites and high-performing algorithms. The listed benchmarks are widely used and recognized within the EMT community, serving as standard testbeds. The recommended algorithms are not arbitrary; they represent the top 5 algorithms based on the Friedman ranking of their average metric performance over 30 independent runs on each benchmark set. The full experimental validation for this ranking is detailed later in Section 5.1. This table serves as an excellent starting point for researchers and can also offer valuable insights when selecting algorithms for new real-world problems. Comprehensive lists, references, and detailed descriptions of all problems, algorithms, and metrics available in MToP are provided in Table LABEL:tab:prob-mt, Table LABEL:tab:algo-mt, and Table 5, respectively.
@@ 4.2.3. Running Experiment and Calculating Metrics
  • For standard multi-task problems (e.g., CEC17-MTSO, CEC19-MTMO), the number of tasks is small, making it practical and recommended to analyze per-task metrics directly, such as Obj, IGD+, or HV. For multi-objective metrics, IGD+ is generally recommended over the traditional IGD because, unlike IGD, it is weakly Pareto-compliant and therefore provides a more comprehensive and stable measure of both convergence and distribution. Furthermore, the HV metric is also provided, as it is a valuable indicator that does not require a known true Pareto front, making it suitable for real-world problems where the optimum is unknown.
@@ 4.2.5. Exporting Experiment Data and Results
  • IOHanalyzer Data (csv): This option exports the complete convergence history (not just the final values) for all selected runs into a csv format. This format is specifically structured for use with the IOHanalyzer (Wang et al., 2022c), facilitating advanced external analysis.
  • Best Decision Variable (mat): This option extracts the decision variables of the best-found solutions of all algorithms and repetitions and saves them to a .mat file for further inspection or use in other applications. For single-objective MTO problems, the best solution is determined by the lowest objective value. For multi-objective MTO problems, the best solutions are a set of Pareto sets (Liu et al., 2024; Choong et al., 2023).
@@ 4.2.6. Visualizing Metric Convergence and Pareto Front
Finally, the module offers integrated plotting tools at the bottom of the results panel (Step 6). By selecting one or more rows and columns in the results table, the user can generate plots. Clicking the Converge button generates a convergence plot for the selected metric results, with a Log toggle available to switch the y-axis to a logarithmic scale, and a Range toggle to display the 95% confidence interval around the mean curve. For multi-objective problems, clicking the Pareto button will plot the final acquired non-dominated solutions (Pareto front approximations). Additionally, the convergence data can be exported in a format compatible with IOHanalyzer (Wang et al., 2022c) for further external analysis and visualization.
@@ 4.4. Executing via Alternative Command Line
A comprehensive set of examples demonstrating this full process including how to instantiate and modify algorithm/problem objects, run experiments, calculate metrics like Obj and IGD+, and programmatically plot the final convergence curves and Pareto front approximations is provided in the cmd_examples.m script located in MToP’s root directory.
@@ 5. Validation and Comparison
This section first performs reproducible validation tests to ensure that the algorithms and problems implemented in MToP are reliable and accurate. Following this, a performance comparison between MToP and a popular EC platform, PlatEMO (Tian et al., 2017), is conducted to demonstrate MToP’s comparable efficiency.
@@ 5.2. Comparison with Other Platforms
Beyond internal validation, it is crucial to benchmark MToP’s performance and efficiency against established, state-of-the-art platforms. We selected PlatEMO (Tian et al., 2017), a widely-used and highly-regarded platform for evolutionary computation, as the basis for our comparison. To ensure a fair and direct comparison, we selected four popular and well-understood single-task multi-objective algorithms that are implemented in both platforms: NSGA-II, SPEA2, MOEA-D, and SMS-EMOA. We executed these algorithms on the standard DTLZ1-7 benchmark suite. Both platforms were run under identical experimental settings, including 30 independent runs and the use of parallel execution.
First, we compared the computational efficiency. Fig. 12 displays the average wall-clock run time for each algorithm across the entire DTLZ suite. The results show that the execution times for MToP and PlatEMO are highly comparable. For NSGA-II, SPEA2, and MOEA-D, the run times are nearly identical. For SMS-EMOA, which involves computationally intensive hypervolume calculations, both platforms show similar overhead. This demonstrates that MToP’s architecture is lightweight and efficient, with no significant computational overhead compared to PlatEMO.
Second, we validated the correctness of our algorithmic implementations by comparing the final solution quality. Table LABEL:tab:results-dtlz presents the mean and standard deviation of the IGD+ metric for 30 runs. A Wilcoxon rank-sum test (at a 0.05 significance level) was conducted between the MToP and PlatEMO results for each case. The + / - / = summary shows that there is no statistically significant difference in the vast majority of cases (24 out of 28). The few minor differences are likely due to small implementation-level variations, but the overall performance is statistically indistinguishable. This similarity in performance is further confirmed visually in Fig. LABEL:fig:pareto-comparison, which plots the final Pareto front approximations from a sample run for each algorithm. The acquired fronts from both platforms are visually coincident.
@@ 6. Discussion and Outlook
Moving forward, the development and enhancement of MToP will continue, drawing inspiration and contributions from the academic community. It is our aspiration that MToP evolves into a valuable tool to propel research in the field of EMT and, by extension, advance the broader evolutionary computation field. Specifically, MToP is considering the incorporation of many-objective optimization codes (Cheng et al., 2016; Deb and Jain, 2014), multitask combinatorial optimization codes (Feng et al., 2021b, a) and more real-world applications (Gupta et al., 2022; Huang et al., 2023) in the future.
@@ Acknowledgements.
This work was partly supported by the National Natural Science Foundation of China (Grant No. 62076225), the Major Research & Development Program of Hubei Province (Grant No. 2025BEB002 & 2023BCA006), the Regional Innovation System Program of Hubei Province (Grant No. 2025EIA022), and the Hubei Superior and Distinctive Discipline Group of “New-Energy Vehicles and Smart Transportation”. All algorithm, problem, and metric implementations within MToP adhere closely to the methodologies outlined in their respective articles. The code provided by the original authors undergoes modifications based on the source code during its integration into MToP. The authors would like to thank the researchers who assisted in modifying the code into the MToP pattern and the researchers who open-sourced their codes. Some of the function implementations in MToP about multi-objective optimization are referred to the code implementation in PlatEMO (Tian et al., 2017), for which the authors are also thankful.